OpenStudy (anonymous):

solve int dx/sinx with x=2t => t := x/2

OpenStudy (anonymous):

\[\int\limits \frac{ dx }{ \sin(x) }\]

OpenStudy (anonymous):

maybe writing \(\sin x\) in terms of \(\tan \frac{x}{2}\) :)

OpenStudy (anonymous):

there's a particular way my professor solves this, i'll show you, but i don't understand the steps

OpenStudy (amistre64):

something like: ln(csc+cot) right?

OpenStudy (anonymous):

the answer is ln abs(sint/cost) + k

OpenStudy (amistre64):

let x = 2t therefore, dx = 2 dt\[\int \frac{dx}{sin(x)}\to \int \frac{2~dt}{sin(2t)}\]

OpenStudy (amistre64):

\[sin(2t)=2sin(t)~cos(t)\] \[\int\frac{dt}{sin(t)~cos(t)}\]

OpenStudy (anonymous):

\[2\int\limits\frac{ dt }{ 2\sin t cost } = \int\limits \frac{ \sin^2t+\cos^2t }{ sint cost }dt= \int\limits tantdt + \int\limits cottdt= -\int\limits \frac{ dcost }{ cost } + \int\limits \frac{ dsint }{ sint }\]

OpenStudy (anonymous):

hmm

OpenStudy (amistre64):

lol, 1 = sin^2 + cos^2 so thats a nice step

OpenStudy (anonymous):

exactly, i understand everything till the - int

OpenStudy (anonymous):

how does int tant dt + int cott dt go to the next step

OpenStudy (anonymous):

been stuck a few hours i think

OpenStudy (amistre64):

\[\int\frac{dt}{sin(t)~cos(t)}\] \[\int\frac{sin^2(t)+cos^2(t)}{sin(t)~cos(t)}dt\] \[\int\frac{sin^2(t)}{sin(t)~cos(t)}+\frac{cos^2(t)}{sin(t)~cos(t)}dt\] \[\int\frac{sin(t)}{cos(t)}+\frac{cos(t)}{sin(t)}dt\] \[\int\frac{sin(t)}{cos(t)}dt+\int\frac{cos(t)}{sin(t)}dt\] \[-ln(cos(t))+ln(sin(t))+k\]

OpenStudy (amistre64):

\[\frac{d}{du}[ln(u)]=\frac{u'}{u}\]

OpenStudy (amistre64):

since sin and cos are derivatives of each other ....

OpenStudy (amistre64):

\[\int\frac{sin(t)}{cos(t)}dt+\int\frac{cos(t)}{sin(t)}dt\] \[\int\frac{-[cos(t)]'}{cos(t)}dt+\int\frac{[sin(t)]'}{sin(t)}dt\] \[-ln(cos(t))+ln(sin(t))+k\]

OpenStudy (anonymous):

alright..let me try and understand

OpenStudy (anonymous):

i don't get it, how did you get to ln?

OpenStudy (amistre64):

its one of the basic derivative rules that you developed from those limit h to 0 stuff

OpenStudy (amistre64):

or if you already know the e setup\[y=ln(u)\] \[e^y=u\] \[y'~e^y = u'\] \[y' = \frac{u'}{e^y}~:~but e^y = u\] \[[ln(u)]' = \frac{u'}{u}\]

OpenStudy (amistre64):

the antiderivative (integral) of the form: u'/u is ln(u)

OpenStudy (anonymous):

i see it now, i seem to have skipped this page

OpenStudy (anonymous):

i'll try again

OpenStudy (amistre64):

:) good luck

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

\[\int\limits\frac{-[\cos(t)]'}{\cos(t)}dt+\int\limits\frac{[\sin(t)]'}{\sin(t)}dt\]

OpenStudy (anonymous):

how'd you get to this?

OpenStudy (anonymous):

because they're derivatives of each other? what do you mean?

OpenStudy (amistre64):

what is the derivative of cos(t)? what is the derivative of sin(t)?

OpenStudy (anonymous):

of cost= sint of sint= -cost

OpenStudy (anonymous):

but why?

OpenStudy (anonymous):

why does that matter?

OpenStudy (amistre64):

you have those a little off [cos(t)]' = -sin(t) [sin(t)]' = cos(t) and it matters because we see that the setups relate to the derivative of the ln(u) function. Which of course is why I went all thru the derivative of ln(u) in the first place .....

OpenStudy (amistre64):

why does this setup solve into ln(...) + k ? because it matters

OpenStudy (amistre64):

@satellite73 any input?

OpenStudy (anonymous):

i get it now

OpenStudy (anonymous):

because - int dcost/cost lets me use int du/u = ln u + k

OpenStudy (anonymous):

which enables me to write it as ln cost

OpenStudy (amistre64):

correct the point is, we have very few formats that we can readily know an integral style to apply. Not all functions can be placed into one of the few formats that can be antiderived. If we can algebra the setup into a familiar format, we can then integrate it up ....

OpenStudy (anonymous):

looks like you are doing fine to me. i have never seen it done like this before

OpenStudy (anonymous):

feeling great right now

OpenStudy (anonymous):

closing this question, and thank you for your help!

OpenStudy (amistre64):

good luck ;)

OpenStudy (anonymous):

btw, i checked and the derivative of cost=sint and sint=-cost